OpenStudy (anonymous):

5) x+2y+3i=3+(2x-y)i 6) 7x+(x-3y)i=3y+9i help me find x and y

OpenStudy (michele_laino):

question 5) hint: please you have to solve this system: \[\begin{gathered} x + 2y = 3 \hfill \\ 3 = 2x - y \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

don't know how elimination works :/ i'm stuck in that part

OpenStudy (michele_laino):

please try to use the substitution method

OpenStudy (michele_laino):

for example from the first equation I get x= 3-2y then I substitute into the second equation, so I can write: 3 = 2(3-2y) - y now please solve that equation for y

OpenStudy (anonymous):

y=3/5?

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

so please substitute y into the equation x = 3 -2y, what do you get?

OpenStudy (michele_laino):

hint: x = 3 - 2*(3/5) = ...?

OpenStudy (anonymous):

x=3/5

OpenStudy (michele_laino):

\[x = 3 - 2 \cdot \frac{3}{5} = 3 - \frac{6}{5} = ...?\]

OpenStudy (michele_laino):

hint: \[x = 3 - 2 \cdot \frac{3}{5} = 3 - \frac{6}{5} = \frac{{15 - 6}}{5} = ...?\]

OpenStudy (anonymous):

9/5

OpenStudy (michele_laino):

perfect!

OpenStudy (michele_laino):

question 6) hint: here, please you have to solve this system: \[\begin{gathered} 7x = 3y \hfill \\ x - 3y = 9 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

wait :)

OpenStudy (anonymous):

@Michele_Laino y=63/18 x=3/2?

OpenStudy (michele_laino):

please I'm checking your answer

OpenStudy (michele_laino):

I got: y= - 7/2, and x= - 3/2